# Thurston’s Teapots and Graph Directed Systems

### v1

@@ 1. Introduction
Figure 1. The horizontal slice at $\lambda=\sqrt{3}$ intersecting with the unit disc.
@@ 2. Graph Directed Iterated Function Systems
One can verify that the definition above is sufficient for running the argument by Solomyak and show that there are asymptotic self similarity around many points. Following the convention in [2, 6], we say a set AA in $\mathbb{C}$ is asymptotically self similar around at point $a$ if there is some $c\in\mathbb{C}$, $|c|>1$, such that the $\epsilon$-ball around 0 of the sets $\{c^{n}(z-a):z\in A\}$ converges under Hausdorff topology. We say two sets AA and $B$ are asymptotically similar to one another around $a\in A$ and $b\in B$, if as $t\to\infty$, the Hausdorff distance between the $\epsilon$ ball around 0 of the set $\{t(z-a):z\in A\}$ and the $\epsilon$ ball around 0 of the set $\{t(z-b):z\in B\}$ converges to 0. Given any infinite path $\gamma$ on $G$ going through consecutive edges $e_{0},e_{1},e_{2},\dots$, denote
@@
As an example, if $\lambda=\frac{\sqrt{5}+1}{2}$ is the golden ratio, then the graph $G$ has 22 vertices $v_{0}$ and $v_{1}$ (with $v_{0}$ being the base vertex), and three edges $e_{1}$, $e_{2}$ and $e_{3}$, such that $e_{1}$ is from $v_{0}$ to $v_{0}$, associated with $(x,z)\mapsto z-zx$, $e_{1}$ is from $v_{0}$ to $v_{1}$, associated with $(x,z)\mapsto zx+2-z$, $e_{2}$ is from $v_{1}$ to $v_{0}$, associated with $(x,z)\mapsto z-zx$.
@@ 3. The Criteria in [4]
Let $\sqrt{2}<\lambda<2$, let $f_{\lambda}$ be the tent map as in Definition $2.$ Let $I_{0}=[0,1-1/\lambda]$, $I_{1}=[1-1/\lambda,1]$. For any $x\in[0,1]$, an itinerary of $x$ under $f_{\lambda}$ is an infinite sequence $It_{\lambda}(x)=(a_{0},a_{1},a_{2},\dots)$, such that $f^{\circ i}(x)\in I_{a_{i}}$. The itinerary of $f_{\lambda}$ is defined as
@@
    If $It_{\lambda}$ starts with $10^{n}1$, there can not be more than $n$ consecutive 11s in $w$.
@@
    When the reverse of the $n$ prefix of $w$ equals the $n$ prefix of $It_{\lambda^{\prime}}$, the number of 11s in this $n$-prefix is odd.
@@ 4. Proof of the Main Theorem
We first construct the parameterized family of graph directed iterated function systems. Consider the finite Markov decomposition of $f_{\lambda}$ formed by splitting interval $[0,1]$ via the forward orbit of the critical point $1-1/\lambda$. Let $\Gamma$ be a directed graph, each vertex corresponding to a Markov block and there is an directed edge from vertex $a$ to vertex $b$ iff the image of the interval corresponding to vertex $a$ under $f_{\lambda}$ contains the interval corresponding to vertex $b$. For every directed edge, if the starting vertex corresponds to a subinterval to the left of $1-1/\lambda$, associate it with the map $(x,z)\mapsto zx+2-z$; if the corresponding subinterval is to the right of $1-1/\lambda$, associate it with the map $(x,z)\mapsto z-zx$. We call the vertex corresponding to the subinterval containing critical value 11 the base vertex $v_{0}$. To construct the family of graph directed iterated function systems, let $G$ be $\Gamma$ with all edges reversed, while keeping the base vertex and all the associated maps.
@@ 4.1. Show that $w$ are all $\lambda$-suitable
From the definition of $\Gamma$, given any finite or infinite path on $\Gamma$, passing through vertices $a_{0},a_{1},\dots,a_{n}$, there is some $x$ in the interval such that $f_{\lambda}^{j}(x)$ is in the subinterval corresponding to vertex $a_{j}$. Hence from the construction of $w$, the reverse of any prefix of $w$ is a section of the itinerary of some point under $f_{\lambda}$. This implies Parts (1) and (2) of Definition 3.5. To show Part (3), suppose a finite word $\alpha$ which is the reverse of a prefix of $w$ is identical to a prefix of some $It_{\lambda^{\prime}}$, then it is also identical to a prefix of $It_{\lambda}$. Now because $w$ comes from a path on $\Gamma$ ending at $v_{0}$, if we concatenate it with the path on $\Gamma$ where the vertices correspond to the subintervals containing $f_{\lambda}^{j}(1)$ (when there is ambiguity pick the subinterval corresponding to the definition $It_{\lambda}$, namely, the subinterval containing $f_{\lambda}^{j}(1-\epsilon)$ for some very small positive $\epsilon$), then we get an infinite path on $\Gamma$. Hence, we can find some $a\in[0,1]$ such that $\alpha$ concatenated with $It_{\lambda}$ is an itinerary in $It_{\lambda}(a)$. If $a<1$, then $\alpha It_{\lambda}<It_{\lambda}$, which implies that the number of 11s in $\alpha$ is odd. If $a=1$, let $\beta_{\lambda}$ be the shortest word such that $It_{\lambda}=\beta_{\lambda}^{\infty}$, $\beta^{\prime}_{\lambda}$ be $\beta_{\lambda}$ with the last letter flipped from 0 to 11 and 11 to 0, then $\alpha It_{\lambda}$ is a concatenation of $\beta_{\lambda}$ or $\beta^{\prime}_{\lambda}$. If $\alpha=\beta^{\prime}_{\lambda}$ then the number of 11s is odd. If $\alpha=\beta_{\lambda}$, suppose $j>0$ is the largest integer such that $It_{\lambda}^{\prime}$ has a prefix of the form $\alpha^{j}$. Remove the common $\alpha^{j}$ prefix of $It_{\lambda}$ and $It_{\lambda}^{\prime}$, we see that $It_{\lambda}^{\prime}$ should have a prefix of the form $\alpha^{j}\gamma$ where $\gamma>\alpha$. Now compare $It_{\lambda}^{\prime}$ and $It_{\lambda}^{\prime}$ with the prefix $\alpha^{j}$ removed, we see that this contradicts with Remark 3.4 hence is impossible.

### v2

@@ 1. Introduction
Figure 1. The horizontal slice at $\lambda=(\sqrt{5}+1)/2$ intersecting with the unit disc.
@@ 2. Graph Directed Iterated Function Systems
One can verify that the definition above is sufficient for running the argument by Solomyak and show that there are asymptotic self similarity around many points. Following the convention in [2, 6], we say a set $A$ in $\mathbb{C}$ is asymptotically self similar around at point $a$ if there is some $c\in\mathbb{C}$, $|c|>1$, such that the $\epsilon$-ball around 0 of the sets $\{c^{n}(z-a):z\in A\}$ converges under Hausdorff topology. We say two sets $A$ and $B$ are asymptotically similar to one another around $a\in A$ and $b\in B$, if as $t\to\infty$, the Hausdorff distance between the $\epsilon$ ball around 0 of the set $\{t(z-a):z\in A\}$ and the $\epsilon$ ball around 0 of the set $\{t(z-b):z\in B\}$ converges to 0. Given any infinite path $\gamma$ on $G$ going through consecutive edges $e_{0},e_{1},e_{2},\dots$, denote
@@
As an example, if $\lambda=\frac{\sqrt{5}+1}{2}$ is the golden ratio, then the graph $G$ has $2$ vertices $v_{0}$ and $v_{1}$ (with $v_{0}$ being the base vertex), and three edges $e_{1}$, $e_{2}$ and $e_{3}$, such that $e_{1}$ is from $v_{0}$ to $v_{0}$, associated with $(x,z)\mapsto z-zx$, $e_{1}$ is from $v_{0}$ to $v_{1}$, associated with $(x,z)\mapsto zx+2-z$, $e_{2}$ is from $v_{1}$ to $v_{0}$, associated with $(x,z)\mapsto z-zx$.
@@ 3. The Criteria in [4]
Let $\sqrt{2}<\lambda<2$, let $f_{\lambda}$ be the tent map as in Definition 2.2. Let $I_{0}=[0,1-1/\lambda]$, $I_{1}=[1-1/\lambda,1]$. For any $x\in[0,1]$, an itinerary of $x$ under $f_{\lambda}$ is an infinite sequence $It_{\lambda}(x)=(a_{0},a_{1},a_{2},\dots)$, such that $f^{\circ i}(x)\in I_{a_{i}}$. The itinerary of $f_{\lambda}$ is defined as
@@
    If $It_{\lambda}$ starts with $10^{n}1$, there can not be more than $n$ consecutive $1$s in $w$.
@@
    When the reverse of the $n$ prefix of $w$ equals the $n$ prefix of $It_{\lambda^{\prime}}$, the number of $1$s in this $n$-prefix is odd.
@@ 4. Proof of the Main Theorem
We first construct the parameterized family of graph directed iterated function systems. Consider the finite Markov decomposition of $f_{\lambda}$ formed by splitting interval $[0,1]$ via the forward orbit of the critical point $1-1/\lambda$. Let $\Gamma$ be a directed graph, each vertex corresponding to a Markov block and there is an directed edge from vertex $a$ to vertex $b$ iff the image of the interval corresponding to vertex $a$ under $f_{\lambda}$ contains the interval corresponding to vertex $b$. For every directed edge, if the starting vertex corresponds to a subinterval to the left of $1-1/\lambda$, associate it with the map $(x,z)\mapsto zx+2-z$; if the corresponding subinterval is to the right of $1-1/\lambda$, associate it with the map $(x,z)\mapsto z-zx$. We call the vertex corresponding to the subinterval containing critical value $1$ the base vertex $v_{0}$. To construct the family of graph directed iterated function systems, let $G$ be $\Gamma$ with all edges reversed, while keeping the base vertex and all the associated maps.
@@ 4.1. Show that $w$ are all $\lambda$-suitable
From the definition of $\Gamma$, given any finite or infinite path on $\Gamma$, passing through vertices $a_{0},a_{1},\dots,a_{n}$, there is some $x$ in the interval such that $f_{\lambda}^{j}(x)$ is in the subinterval corresponding to vertex $a_{j}$. Hence from the construction of $w$, the reverse of any prefix of $w$ is a section of the itinerary of some point under $f_{\lambda}$. This implies Parts (1) and (2) of Definition 3.5. To show Part (3), suppose a finite word $\alpha$ which is the reverse of a prefix of $w$ is identical to a prefix of some $It_{\lambda^{\prime}}$, then it is also identical to a prefix of $It_{\lambda}$. Now because $w$ comes from a path on $\Gamma$ ending at $v_{0}$, if we concatenate it with the path on $\Gamma$ where the vertices correspond to the subintervals containing $f_{\lambda}^{j}(1)$ (when there is ambiguity pick the subinterval corresponding to the definition $It_{\lambda}$, namely, the subinterval containing $f_{\lambda}^{j}(1-\epsilon)$ for some very small positive $\epsilon$), then we get an infinite path on $\Gamma$. Hence, we can find some $a\in[0,1]$ such that $\alpha$ concatenated with $It_{\lambda}$ is an itinerary in $It_{\lambda}(a)$. If $a<1$, then $\alpha It_{\lambda}<It_{\lambda}$, which implies that the number of $1$s in $\alpha$ is odd. If $a=1$, let $\beta_{\lambda}$ be the shortest word such that $It_{\lambda}=\beta_{\lambda}^{\infty}$, $\beta^{\prime}_{\lambda}$ be $\beta_{\lambda}$ with the last letter flipped from 0 to $1$ and $1$ to 0, then $\alpha It_{\lambda}$ is a concatenation of $\beta_{\lambda}$ or $\beta^{\prime}_{\lambda}$. If $\alpha=\beta^{\prime}_{\lambda}$ then the number of $1$s is odd. If $\alpha=\beta_{\lambda}$, suppose $j>0$ is the largest integer such that $It_{\lambda}^{\prime}$ has a prefix of the form $\alpha^{j}$. Remove the common $\alpha^{j}$ prefix of $It_{\lambda}$ and $It_{\lambda}^{\prime}$, we see that $It_{\lambda}^{\prime}$ should have a prefix of the form $\alpha^{j}\gamma$ where $\gamma>\alpha$. Now compare $It_{\lambda}^{\prime}$ and $It_{\lambda}^{\prime}$ with the prefix $\alpha^{j}$ removed, we see that this contradicts with Remark 3.4 hence is impossible.